countryboy:

What is the approximate volume of a cone with a height of 6 mm and radius of 18 mm? Use 3.14 to approximate pi, and express your final answer to the nearest hundredth. Enter your answer in the box. mm³

Nnesha:

\[\large\rm V=\frac{1}{3}\pi r^2 h\] V=volume pi=3.14 h= 6mm r=18mm plugin the values and simplify.